THE LIFE OF ON. THEİ MORTGAGE BEGAN İDEC.L Hous mwan WAS PAID ONTHE MORTGAde DuRiNC NTEREs
Answer #1
Pv Amount of loan $525,000
Rate Monthly interest rate=(6.6/12)% 0.55%
Nper Number of months of mortgage 336 (28years*12)
PMT Monthly Payment $3,430.77 (Using PMT function of excelwith Rate=0.55%,Nper=336,Pv=-525000)
INTEREST PAID IN FIRST CALENDAR YEAR
A B C=A*0.55% D=B-C E=A-D
Payment on Month Beginning Balance TotalPayment Interest Principal Ending Balance
January,1 1 $525,000 $3,430.77 $2,887.50 $543.27 $524,456.73
February, 1 2 $524,456.73 $3,430.77 $2,884.51 $546.26 $523,910.48
March,1 3 $523,910.48 $3,430.77 $2,881.51 $549.26 $523,361.22
April,1 4 $523,361.22 $3,430.77 $2,878.49 $552.28 $522,808.93
May,1 5 $522,808.93 $3,430.77 $2,875.45 $555.32 $522,253.62
June,1 6 $522,253.62 $3,430.77 $2,872.39 $558.37 $521,695.24
July,1 7 $521,695.24 $3,430.77 $2,869.32 $561.44 $521,133.80
August,1 8 $521,133.80 $3,430.77 $2,866.24 $564.53 $520,569.27
Sep,1 9 $520,569.27 $3,430.77 $2,863.13 $567.64 $520,001.63
Oct,1 10 $520,001.63 $3,430.77 $2,860.01 $570.76 $519,430.87
Nov,1 11 $519,430.87 $3,430.77 $2,856.87 $573.90 $518,856.97
Dec,1 12 $518,856.97 $3,430.77 $2,853.71 $577.05 $518,279.92
TOTAL $41,169.22 $34,449.13 $6,720.08
Interest paid in first calendar year $34,449.13